Question Three

  1. Explain the effects of inadequate working capital to a firm
  2. An investor has an investment capital of Sh.2,000,000. He wishes to invest in two securities, A and B in the following proportion; Sh.400,000 in security A and Sh.1, 600,000 in security B. The returns on these two securities depend on the state of the economy as shown below:

State of Economy

Probability

Return on Security A

Return on security B

Boom

0.4

18%

24%

Normal

0.5

14%

22%

Recession

0.1

12%

21%

  1. Compute the expected portfolio return
  2. Determine the correlation coefficient between security A and Security B and interpret it.
  3. Calculate the portfolio risk as measured by standard deviation.
Answer #1
Total Portfolio value = Value of Sec A + Value of Sec B
=400000+1600000
=2000000
Weight of Sec A = Value of Sec A/Total Portfolio Value
= 400000/2000000
=0.2
Weight of Sec B = Value of Sec B/Total Portfolio Value
= 1600000/2000000
=0.8

i

Sec A
Scenario Probability Return% =rate of return% * probability Actual return -expected return(A)% (A)^2* probability
Boom 0.4 18 7.2 2.6 0.0002704
Normal 0.5 14 7 -1.4 9.8E-05
Recession 0.1 12 1.2 -3.4 0.0001156
Expected return %= sum of weighted return = 15.4 Sum=Variance Sec A= 0.00048
Standard deviation of Sec A% =(Variance)^(1/2) 2.2
Sec B
Scenario Probability Return% =rate of return% * probability Actual return -expected return(A)% (B)^2* probability
Boom 0.4 24 9.6 1.3 6.76E-05
Normal 0.5 22 11 -0.7 2.45E-05
Recession 0.1 21 2.1 -1.7 2.89E-05
Expected return %= sum of weighted return = 22.7 Sum=Variance Sec B= 0.00012
Standard deviation of Sec B% =(Variance)^(1/2) 1.1

i

Expected return%= Wt Sec A*Return Sec A+Wt Sec B*Return Sec B
Expected return%= 0.2*15.4+0.8*22.7
Expected return%= 21.24

ii

Covariance Sec A Sec B:
Scenario Probability Actual return% -expected return% for A(A) Actual return% -expected return% For B(B) (A)*(B)*probability
Boom 0.4 2.6 1.3 0.0001352
Normal 0.5 -1.4 -0.7 4.9E-05
Recession 0.1 -3.4 -1.7 5.78E-05
Covariance=sum= 0.000242
Correlation A&B= Covariance/(std devA*std devB)= 1

iii

Variance =( w2A*σ2(RA) + w2B*σ2(RB) + 2*(wA)*(wB)*Cor(RA, RB)*σ(RA)*σ(RB))
Variance =0.2^2*0.022^2+0.8^2*0.011^2+2*0.2*0.8*0.022*0.011*1
Variance 0.000170
Standard deviation= (variance)^0.5
Standard deviation= 1.30%